# Flow to the Mode: Mode-Seeking Diffusion Autoencoders for State-of-the-Art Image Tokenization

### v1

@@ Abstract
Since the advent of popular visual generation frameworks like VQGAN and latent diffusion models, state-of-the-art image generation systems have generally been two-stage systems that first tokenize or compress visual data into a lower-dimensional latent space before learning a generative model. Tokenizer training typically follows a standard recipe in which images are compressed and reconstructed subject to a combination of MSE, perceptual, and adversarial losses. Diffusion autoencoders have been proposed in prior work as a way to learn end-to-end perceptually-oriented image compression, but have not yet shown state-of-the-art performance on the competitive task of ImageNet-1K reconstruction. We propose FlowMo, a transformer-based diffusion autoencoder that achieves a new state-of-the-art for image tokenization at multiple compression rates without using convolutions, adversarial losses, spatially-aligned two-dimensional latent codes, or distilling from other tokenizers. Our key insight is that FlowMo training should be broken into a mode-matching pre-training stage and a mode-seeking post-training stage. In addition, we conduct extensive analyses and explore the training of generative models atop the FlowMo tokenizer. Our code and model will be available at https://kylesargent.github.io/flowmo.
@@ 1 Introduction
Generative models such as diffusion models [49, 50, 24, 17, 36, 35] and discrete autoregressive models [16, 8] have seen compelling applications in the creation of image and video content [6, 47, 32]. Although less common systems such as cascaded [47] and single-stage [21, 25] systems have been explored, state-of-the-art systems for visual generation generally consist of two stages: first, a “tokenizer” is learned to compress pixel data to a smaller and more tractable discrete or continuous latent space; second, a generative model is trained over this compressed latent space. Due to the dominance of this two-stage paradigm, tokenization has become an active research area in its own right, with multiple influential works focusing heavily on tokenizer training and design [61, 37, 56, 60, 18]. In this work, we focus on tokenizers with a discrete latent space.
@@
Since VQGAN [16], a dominant architecture and training scheme for discrete image tokenizers has emerged. State-of-the-art image tokenizers are typically convolutional autoencoders trained to downsample visual data to a 2-dimensional, spatially-aligned latent code, and then upsample the latent code to a reconstruction regularized by adversarial and perceptual losses. Various deviations from this setup have been proposed: TiTok [61] used a transformer-based architecture and 1-dimensional latent code although relying on a traditional CNN-based tokenizer for an initial distillation stage, and ViT-VQGAN [59] used a transformer-based encoder and decoder. In general, however, the same basic setup for image tokenization has remained dominant.
In this work, we propose FlowMo (Flow to the Mode), a significant technical departure from the current state-of-the-art for image tokenization. First, in tackling the task of image tokenization, we propose to model the multimodal reconstruction distribution using a rectified flow objective for the decoder [36, 35]. Second, we use a fully transformer-based architecture to encode to and decode from a 1-dimensional latent code. Third, we optimize FlowMo end-to-end, without distilling from a pre-existing 2-dimensional tokenizer [61] or encoding on top of a latent space from a 2-dimensional tokenizer [1].
Most importantly, FlowMo achieves state-of-the-art performance via a key insight. FlowMo is a diffusion autoencoder, a system which has been explored in prior work [45, 2, 26, 58]. Yet, the state-of-the-art for the most competitive perceptual reconstruction benchmark, ImageNet-1K [13], continues to be dominated by more traditional CNN- or GAN-based tokenizers. FlowMo achieves state-of-the-art tokenization via the following key insight: for the task of perceptual reconstruction, it is better to sample modes of the reconstruction distribution that are perceptually close to the original image than to try to match all modes. Hence, we divide FlowMo training into a mode-matching pre-training stage, in which the system is trained end-to-end with a diffusion loss on the decoder like typical diffusion autoencoders, and a mode-seeking post-training stage, in which the system is trained to selectively drop modes of the reconstruction distribution which are not close to the original image. We explain both stages in Section 3.
@@
  • We present a simple but novel architecture for image tokenization based on diffusion autoencoders [45] and the multimodal diffusion image transformer (MMDiT) [17].
@@ 2 Related Work
Image tokenization. State-of-the-art systems for visual generation generally consist of two stages (with a few notable exceptions attempting to learn directly in pixel space [25, 27, 21, 12]). The first stage is a “tokenizer” which reduces the spatial or spatiotemporal dimension of pixel data by projecting to a continuous or discrete latent space. The dominance of this paradigm has led to numerous works that study tokenizers as important components in their own right [9, 61, 56, 60, 37, 52]. In this work, we study tokenizers with discrete latents. Unlike prior work, FlowMo is the first diffusion autoencoder-based architecture to achieve state-of-the-art performance on ImageNet-1K reconstruction.
Diffusion autoencoders. Diffusion models are popular for visual generation [46, 6], and simplified frameworks such as rectified flow [35, 36] have led to further adoption. The idea of learning an autoencoder end-to-end with a diffusion decoder was first proposed in [45]. Various works have followed up on this idea, studying diffusion autoencoders for representation learning [28] and in particular for perceptually-oriented image compression [58, 2, 41].
Diffusion post-training. Various work has attempted to improve the quality of diffusion model samples with dedicated post-training strategies, the goal of which is generally to instill desired attributes such as aesthetic quality. DDPO [3] and DPOK [19] explore reinforcement learning objectives, while DRAFT [11] and AlignProp [44] explore backpropagation through the sampling chain. FlowMo uses a backpropagation-based post-training strategy, adapted to the continuous noise schedule of rectified flow and the unique setting of image tokenization.
Concurrent work. In work concurrent to ours, DiTo [10] studies diffusion autoencoders for the continuous image tokenization task. Unlike DiTo, we focus on discrete image tokenization and compare against the strongest state-of-the-art discrete tokenizers available [37, 52]. Meanwhile, FlexTok [1] proposes a system in which diffusion tokenizers are learned on top of a latent space from a traditional continuous variational autoencoder (VAE) trained with perceptual, adversarial, and reconstruction losses. FlowMo does not rely on an auxiliary VAE. The reader may consult these works for a more complete picture of this developing field.
@@ 3 Method
Existing state-of-the-art tokenizers have enabled considerable progress in visual generation. However, they have a variety of drawbacks. First, they require adversarial losses [60, 37, 16], which can be unstable and difficult to tune. They almost all require CNNs in at least one training stage, making it difficult to leverage the hardware efficiency and well-understood scaling behavior of transformers [54, 57]. Finally, they often leverage distillation from previously trained tokenizers [1, 61] in order to achieve state-of-the-art performance.
@@
  • Diffusion-based. Prior discrete tokenizers leverage adversarial losses to sample a plausible mode of the reconstruction distribution, which necessitates adaptive gradient scale computation [16], LeCam regularization [55, 60], or careful loss weight tuning for stability. Instead, we will use a diffusion decoder, given that diffusion models have proven simple, well-suited for multi-modal modeling, and reliable at scale [6, 17].
  • Transformer-only, with 1-dimensional latent code. Nearly every state-of-the-art image tokenization framework uses a CNN-based architecture or predicts a spatially aligned 2-dimensional latent code, with the exception of TiTok [61], which still relies on distilling from a pre-trained CNN-based tokenizer with a 2-dimensional latent code. While these choices may provide a useful inductive bias, we feel transformer-based tokenizers with 1-dimensional latent codes may eventually provide more efficiency and flexibility at large data and model scales. FlowMo learns 1-dimensional latent codes in a transformer-based architecture following MMDiT.
  • State-of-the-art. There has already been considerable work in diffusion autoencoders [45, 26, 2, 28], but FlowMo is the first to achieve state-of-the-art perceptual reconstruction on ImageNet-1K.
@@
We will now describe FlowMo in detail. First, we will go over the FlowMo architecture, which is a simple transformer-based architecture based on MMDiT [17]. Then, we will explain the two training stages: mode-matching pre-training (Stage 1A) and mode-seeking fine-tuning (Stage 1B). Finally, we will discuss generative modeling over the FlowMo latent space (Stage 2) and sampling.
@@ 3.1 Architecture
FlowMo consists of an encoder $e_{\theta}$ and decoder $d_{\theta}$. Both are transformers based on the multimodal diffusion image transformer (MMDiT) [17]. Given a patchified [15] image $x\in\mathbb{R}^{n}$ and an initial latent code $c_{0}\in\mathbb{R}^{d}$ (a vector of all zeroes), the encoder produces a latent token sequence
@@
by interacting the concatenated sequence of latent and image tokens with self-attention while maintaining separate streams for each modality. $\hat{c}$ is then binarized elementwise via a quantization operation $q$ following lookup-free quantization (LFQ) [60], yielding
@@
Following rectified flow [36], the decoder is trained to model a velocity field $v$ from noise to data defined as
@@
The decoder processes $x_{t}$ and $c$ identically to how the encoder processes $x$ and $c_{0}$, but additionally accepts the time (or noise level) parameter $t$ to condition each MMDiT block via AdaLN modulation [43]. The encoder $e_{\theta}$ and decoder $d_{\theta}$ are architecturally symmetric but differently sized, with the decoder being considerably larger and deeper. We use the $\mu P$ parameterization [57] for all models in order to simplify hyperparameter transfer between exploratory and scaled-up configurations.
@@ 3.2 Stage 1A: Mode-matching pre-training
In the first training stage, our goal is to train the encoder $e_{\theta}$ and decoder $d_{\theta}$ jointly so that the quantized $c$ is maximally informative about $x$, and so that $p_{\theta}(x|c)$ matches the true distribution $p(x|c)$, which is necessarily multimodal since the quantized $c$ contains only limited information. A full diagram of Stage 1A is given in Figure 4. For theoretical justification, we appeal to prior work in diffusion autoencoders [58], which notes that the end-to-end diffusion objective corresponds to a modified variational lower bound with non-Gaussian decoder $p(x|c)$.
FlowMo is trained end-to-end as a diffusion autoencoder [45]. Specifically, the FlowMo encoder and decoder are trained end-to-end to optimize the rectified flow loss $\mathcal{L}_{\text{flow}}$ on the decoder output. Given noise $z\,{\sim}\,\mathcal{N}(0,I)$, data $x\,{\sim}\,p_{x}$, and time (or noise level) $t\,{\sim}\,p_{t}$, $t\in[0,1]$, we define
@@
We also use a learned perceptual distance $d_{\text{perc}}$ [29] to supervise the 1-step denoised prediction of the network via
@@
Finally, on the latent code $c$ we use the entropy and commitment losses of LFQ following [60], with
$\mathcal{L}_{\text{ent}}=\mathbb{E}\big{[}H(q(\hat{c})\big{]}-H(\mathbb{E}\big{[}q(\hat{c})\big{]}),$ (7)
@@
For simplicity, these two losses could be replaced with finite scalar quantization (FSQ) [39], but we find LFQ to perform slightly better. Our training loss in Stage 1A is
@@
with loss weights and further details given in the supplementary material. Images are rescaled to lie in $[-1,1]$. The noise level $t$ is sampled from a thick-tailed logit-normal distribution following Stable Diffusion 3 [17].
@@ 3.3 Stage 1B: Mode-seeking post-training
In this stage, our goal is to optimize the decoder distribution $p_{\theta}(x|c)$ to seek modes which are perceptually of high similarity to the original image. We accomplish this by freezing the encoder and co-training the decoder on $\mathcal{L}_{\text{flow}}$ along with a post-training objective $\mathcal{L}_{\text{sample}}$ inspired by prior work on diffusion model post-training [11, 44].
@@
$\mathcal{L}_{\text{sample}}=\mathbb{E}\bigg{[}d_{\text{perc}}(x,d_{t_{n}}\circ d_{t_{n-1}}\circ\cdots\circ d_{t_{1}}(x)\bigg{]},$ (11)
@@
We find the value of $\lambda_{\text{sample}}$ in this stage to be particularly important, and use $\lambda_{\text{sample}}=0.01$. Too small a weight on $\mathcal{L}_{\text{sample}}$ results in poor rFID as the network forgets the perceptual features acquired in Stage 1A, whereas too large a weight results in either so-called “reward-hacking” as the decoder $d_{\theta}$ overfits to $d_{\text{perc}}$, or training divergence. We use an LPIPS-VGG network [62] as the perceptual network in Stage 1A, but, following [56], we use a ResNet as the perceptual network in Stage 1B. This and other design choices are explained in the supplementary material, and we justify there via ablation that simply using the ResNet network on the 1-step denoised prediction as in $\mathcal{L}_{perc}$ is ineffective; the post-training stage presented is necessary. This stage is computationally expensive as it requires backpropagation through the full sampling chain. We use $n=8$ and find it generalizes well to other $n$ at sampling time.
@@
0.006 TiTok-L-32 [61] 32 $2^{12}$ 2.21 15.60 0.359 0.322
0.012 TiTok-B-64 [61] 64 $2^{12}$ 1.70 16.80 0.407 0.252
0.023 TiTok-S-128 [61] 128 $2^{12}$ 1.71 17.52 0.437 0.210
0.055 LlamaGen-16 [52] 256 $2^{14}$ 2.19 20.67 0.589 0.132
MaskBiT [56] 256 $2^{14}$ 1.37 21.5 0.56 -
0.062 Cosmos DI-16x16 [61] 256 $\approx 2^{16}$ 4.40 19.98 0.536 0.153
0.070 OpenMagViT-V2 [37] 256 $2^{18}$ 1.17 21.63 0.640 0.111
@@
0.203 ViT-VQGAN [59] 1024 $2^{13}$ 1.28 - - -
0.219 LlamaGen-32 [52] 1024 $2^{14}$ 0.59 24.44 0.768 0.064

### v3

@@ Abstract
Since the advent of popular visual generation frameworks like VQGAN and latent diffusion models, state-of-the-art image generation systems have generally been two-stage systems that first tokenize or compress visual data into a lower-dimensional latent space before learning a generative model. Tokenizer training typically follows a standard recipe in which images are compressed and reconstructed subject to a combination of MSE, perceptual, and adversarial losses. Diffusion autoencoders have been proposed in prior work as a way to learn end-to-end perceptually-oriented image compression, but have not yet shown state-of-the-art performance on the competitive task of ImageNet-1K reconstruction. We propose FlowMo, a transformer-based diffusion autoencoder that achieves a new state-of-the-art for image tokenization at multiple compression rates without using convolutions, adversarial losses, spatially-aligned two-dimensional latent codes, or distilling from other tokenizers. Our key insight is that FlowMo training should be broken into a mode-matching pre-training stage and a mode-seeking post-training stage. In addition, we conduct extensive analyses and explore the training of generative models atop the FlowMo tokenizer. Our code and models are available at https://kylesargent.github.io/flowmo.
@@ 1 Introduction
Generative models such as diffusion models [diffusion2015, yangsonggradients, ddpm, sd3, rectified_flow, flow_matching] and discrete autoregressive models [vqgan, maskgit] have seen compelling applications in the creation of image and video content [sora, imagen, videpoet]. Although less common systems such as cascaded [imagen] and single-stage [matryoshka, simple_diffusion] systems have been explored, state-of-the-art systems for visual generation generally consist of two stages: first, a “tokenizer” is learned to compress pixel data to a smaller and more tractable discrete or continuous latent space; second, a generative model is trained over this compressed latent space. Due to the dominance of this two-stage paradigm, tokenization has become an active research area in its own right, with multiple influential works focusing heavily on tokenizer training and design [titok, openmagvit_v2, maskbit, magvit_v2, cosmos]. In this work, we focus on tokenizers with a discrete latent space.
@@
Since VQGAN [vqgan], a dominant architecture and training scheme for discrete image tokenizers has emerged. State-of-the-art image tokenizers are typically convolutional autoencoders trained to downsample visual data to a 2-dimensional, spatially-aligned latent code, and then upsample the latent code to a reconstruction regularized by adversarial and perceptual losses. Various deviations from this setup have been proposed: TiTok [titok] used a transformer-based architecture and 1-dimensional latent code although relying on a traditional CNN-based tokenizer for an initial distillation stage, and ViT-VQGAN [vitvqgan] used a transformer-based encoder and decoder. In general, however, the same basic setup for image tokenization has remained dominant.
In this work, we propose FlowMo (Flow to the Mode), a significant technical departure from the current state-of-the-art for image tokenization. First, in tackling the task of image tokenization, we propose to model the multimodal reconstruction distribution using a rectified flow objective for the decoder [rectified_flow, flow_matching]. Second, we use a fully transformer-based architecture to encode to and decode from a 1-dimensional latent code. Third, we optimize FlowMo end-to-end, without distilling from a pre-existing 2-dimensional tokenizer [titok] or encoding on top of a latent space from a 2-dimensional tokenizer [flextok].
Most importantly, FlowMo achieves state-of-the-art performance via a key insight. FlowMo is a diffusion autoencoder, a system which has been explored in prior work [diffae, swycc, highfidelity, yangmandt]. Yet, the state-of-the-art for the most competitive perceptual reconstruction benchmark, ImageNet-1K [imagenet], continues to be dominated by more traditional CNN- or GAN-based tokenizers. FlowMo achieves state-of-the-art tokenization via the following key insight: for the task of perceptual reconstruction, it is better to sample modes of the reconstruction distribution that are perceptually close to the original image than to try to match all modes. Hence, we divide FlowMo training into a mode-matching pre-training stage, in which the system is trained end-to-end with a diffusion loss on the decoder like typical diffusion autoencoders, and a mode-seeking post-training stage, in which the system is trained to selectively drop modes of the reconstruction distribution which are not close to the original image. We explain both stages in Section 3.
@@
  • We present a simple but novel architecture for image tokenization based on diffusion autoencoders [diffae] and the multimodal diffusion image transformer (MMDiT) [sd3].
@@ 2 Related Work
Image tokenization. State-of-the-art systems for visual generation generally consist of two stages (with a few notable exceptions attempting to learn directly in pixel space [simple_diffusion, simpler_diffusion, matryoshka, hourglass]). The first stage is a “tokenizer” which reduces the spatial or spatiotemporal dimension of pixel data by projecting to a continuous or discrete latent space. The dominance of this paradigm has led to numerous works that study tokenizers as important components in their own right [dcae, titok, maskbit, magvit_v2, openmagvit_v2, llamagen]. In this work, we study tokenizers with discrete latents. Unlike prior work, FlowMo is the first diffusion autoencoder-based architecture to achieve state-of-the-art performance on ImageNet-1K reconstruction.
Diffusion autoencoders. Diffusion models are popular for visual generation [ldm, sora], and simplified frameworks such as rectified flow [flow_matching, rectified_flow] have led to further adoption. The idea of learning an autoencoder end-to-end with a diffusion decoder was first proposed in [diffae]. Various works have followed up on this idea, studying diffusion autoencoders for representation learning [soda] and in particular for perceptually-oriented image compression [yangmandt, swycc, dalle3]. $\epsilon$-VAE [evae] explores a diffusion autoencoder objective in the context of GAN-based tokenizer trained with adversarial and perceptual losses, while DiscoDiff [discodiff] uses a diffusion autoencoder to learning high-level semantic discrete latents for images prior to an autoregressive generation stage.
Diffusion post-training. Various work has attempted to improve the quality of diffusion model samples with dedicated post-training strategies, the goal of which is generally to instill desired attributes such as aesthetic quality. DDPO [ddpo] and DPOK [dpok] explore reinforcement learning objectives, while DRAFT [draft] and AlignProp [alignprop] explore backpropagation through the sampling chain. FlowMo uses a backpropagation-based post-training strategy, adapted to the continuous noise schedule of rectified flow and the unique setting of image tokenization.
Concurrent work. In work concurrent to ours, DiTo [dito] studies diffusion autoencoders for the continuous image tokenization task. Unlike DiTo, we focus on discrete image tokenization and compare against the strongest state-of-the-art discrete tokenizers available [openmagvit_v2, llamagen]. Meanwhile, FlexTok [flextok] proposes a system in which diffusion tokenizers are learned on top of a latent space from a traditional continuous variational autoencoder (VAE) trained with perceptual, adversarial, and reconstruction losses. FlowMo does not rely on an auxiliary VAE. The reader may consult these works for a more complete picture of this developing field.
@@ 3 Method
Existing state-of-the-art tokenizers have enabled considerable progress in visual generation. However, they have a variety of drawbacks. First, they require adversarial losses [magvit_v2, openmagvit_v2, vqgan], which can be unstable and difficult to tune. They almost all require CNNs in at least one training stage, making it difficult to leverage the hardware efficiency and well-understood scaling behavior of transformers [llama, mup]. Finally, they often leverage distillation from previously trained tokenizers [flextok, titok] in order to achieve state-of-the-art performance.
@@
  • Diffusion-based. Prior discrete tokenizers leverage adversarial losses to sample a plausible mode of the reconstruction distribution, which necessitates adaptive gradient scale computation [vqgan], LeCam regularization [lecam, magvit_v2], or careful loss weight tuning for stability. Instead, we will use a diffusion decoder, given that diffusion models have proven simple, well-suited for multi-modal modeling, and reliable at scale [sora, sd3].
  • Transformer-only, with 1-dimensional latent code. Nearly every state-of-the-art image tokenization framework uses a CNN-based architecture or predicts a spatially aligned 2-dimensional latent code, with the exception of TiTok [titok], which still relies on distilling from a pre-trained CNN-based tokenizer with a 2-dimensional latent code. While these choices may provide a useful inductive bias, we feel transformer-based tokenizers with 1-dimensional latent codes may eventually provide more efficiency and flexibility at large data and model scales. FlowMo learns 1-dimensional latent codes in a transformer-based architecture following MMDiT.
  • State-of-the-art. There has already been considerable work in diffusion autoencoders [diffae, highfidelity, swycc, soda], but FlowMo is the first to achieve state-of-the-art perceptual reconstruction on ImageNet-1K.
@@
We will now describe FlowMo in detail. First, we will go over the FlowMo architecture, which is a simple transformer-based architecture based on MMDiT [sd3]. Then, we will explain the two training stages: mode-matching pre-training (Stage 1A) and mode-seeking fine-tuning (Stage 1B). Finally, we will discuss generative modeling over the FlowMo latent space (Stage 2) and sampling.
@@ 3.1 Architecture
FlowMo consists of an encoder $e_{\theta}$ and decoder $d_{\theta}$. Both are transformers based on the multimodal diffusion image transformer (MMDiT) [sd3] adapted from Flux [flux2024]. Given a patchified [vit] image $x\in\mathbb{R}^{n}$ and an initial latent code $c_{0}\in\mathbb{R}^{d}$ (a vector of all zeroes), the encoder produces a latent token sequence
@@
by interacting the concatenated sequence of latent and image tokens with self-attention while maintaining separate streams for each modality. $\hat{c}$ is then binarized elementwise via a quantization operation $q$ following lookup-free quantization (LFQ) [magvit_v2], yielding
@@
Following rectified flow [rectified_flow], the decoder is trained to model a velocity field $v$ from noise to data defined as
@@
The decoder processes $x_{t}$ and $c$ identically to how the encoder processes $x$ and $c_{0}$, but additionally accepts the time (or noise level) parameter $t$ to condition each MMDiT block via AdaLN modulation [DiT]. The encoder $e_{\theta}$ and decoder $d_{\theta}$ are architecturally symmetric but differently sized, with the decoder being considerably larger and deeper. We use the $\mu P$ parameterization [mup] for all models in order to simplify hyperparameter transfer between exploratory and scaled-up configurations.
@@ 3.2 Stage 1A: Mode-matching pre-training
In the first training stage, our goal is to train the encoder $e_{\theta}$ and decoder $d_{\theta}$ jointly so that the quantized $c$ is maximally informative about $x$, and so that $p_{\theta}(x|c)$ matches the true distribution $p(x|c)$, which is necessarily multimodal since the quantized $c$ contains only limited information. A full diagram of Stage 1A is given in Figure 4. For theoretical justification, we appeal to prior work in diffusion autoencoders [yangmandt], which notes that the end-to-end diffusion objective corresponds to a modified variational lower bound with non-Gaussian decoder $p(x|c)$.
FlowMo is trained end-to-end as a diffusion autoencoder [diffae]. Specifically, the FlowMo encoder and decoder are trained end-to-end to optimize the rectified flow loss $\mathcal{L}_{\text{flow}}$ on the decoder output. Given noise $z\,{\sim}\,\mathcal{N}(0,I)$, data $x\,{\sim}\,p_{x}$, and time (or noise level) $t\,{\sim}\,p_{t}$, $t\in[0,1]$, we define
@@
We also use a learned perceptual distance $d_{\text{perc}}$ [perceptual_loss] to supervise the 1-step denoised prediction of the network via
@@
Finally, on the latent code $c$ we use the entropy and commitment losses of LFQ following [magvit_v2], with
[…]
@@
For simplicity, these two losses could be replaced with finite scalar quantization (FSQ) [fsq], but we find LFQ to perform slightly better. Our training loss in Stage 1A is
@@
with loss weights and further details given in the supplementary material. Images are rescaled to lie in [1,1][-1,1]. The noise level $t$ is sampled from a thick-tailed logit-normal distribution following Stable Diffusion 3 [sd3].
@@ 3.3 Stage 1B: Mode-seeking post-training
In this stage, our goal is to optimize the decoder distribution $p_{\theta}(x|c)$ to seek modes which are perceptually of high similarity to the original image. We accomplish this by freezing the encoder and co-training the decoder on $\mathcal{L}_{\text{flow}}$ along with a post-training objective $\mathcal{L}_{\text{sample}}$ inspired by prior work on diffusion model post-training [draft, alignprop].
@@
$\mathcal{L}_{\text{sample}}=\mathbb{E}\bigg[d_{\text{perc}}\left(x,d_{t_{n}}\circ d_{t_{n-1}}\circ\cdots\circ d_{t_{1}}(z)\right)\bigg],$ (11)
@@
We find the value of $\lambda_{\text{sample}}$ in this stage to be particularly important, and use $\lambda_{\text{sample}}=0.01$. Too small a weight on $\mathcal{L}_{\text{sample}}$ results in poor rFID as the network forgets the perceptual features acquired in Stage 1A, whereas too large a weight results in either so-called “reward-hacking” as the decoder $d_{\theta}$ overfits to $d_{\text{perc}}$, or training divergence. We use an LPIPS-VGG network [lpips] as the perceptual network in Stage 1A, but, following [maskbit], we use a ResNet as the perceptual network in Stage 1B. This and other design choices are explained in the supplementary material, and we justify there via ablation that simply using the ResNet network on the 1-step denoised prediction as in $\mathcal{L}_{perc}$ is ineffective; the post-training stage presented is necessary. This stage is computationally expensive as it requires backpropagation through the full sampling chain. We use $n=8$ and find it generalizes well to other $n$ at sampling time.
@@
0.006 TiTok-L-32 [titok] 32 $^{12} 2.2$1 15.60 0.359 0.322
0.012 TiTok-B-64 [titok] 64 $2^{12}$ 1.70 16.80 0.407 0.252
0.023 TiTok-S-128 [titok] 128 $2^{12}$ 1.71 17.52 0.437 0.210
0.055 LlamaGen-16 [llamagen] 256 $2^{14}$ 2.19 20.67 0.589 0.132
MaskBiT [maskbit] 256 $2^{14}$ 1.37 21.5 0.56 -
0.062 Cosmos DI-16x16 [titok] 256 $\approx 2^{16}$ 4.40 19.98 0.536 0.153
0.070 OpenMagViT-V2 [openmagvit_v2] 256 $2^{18}$ 1.17 21.63 0.640 0.111
@@
0.203 ViT-VQGAN [vitvqgan] 1024 $2^{13}$ 1.28 - - -
0.219 LlamaGen-32 [llamagen] 1024 $2^{14}$ 0.59 24.44 0.768 0.064